# Deflanderization for Game Dialogue: Balancing Character Authenticity with Task Execution in LLM-based NPCs

### v1

@@ Abstract
Pasin Buakhaw1, Kun Kerdthaisong2, Phuree Phenhiran2,†, Pitikorn Khlaisamniang3, Supasate Vorathammathorn3, Piyalitt Ittichaiwong4,5,*, Nutchanon Yongsatianchot2,* 1 Department of Computer Engineering and Digital Technology, Faculty of Engineering, Chulalongkorn University 2Thammasat School of Engineering, Thammasat University 3Artificial Intelligence Association of Thailand 4School of Biomedical Engineering & Imaging Sciences, King’s College London 5Siriraj Informatics and Data Innovation Center (SIData+), Faculty of Medicine, Siriraj Hospital, Mahidol University These authors contributed equally to this work. *Corresponding author
@@ 1 Introduction
First, user persona-LLMs are designed as purpose-built assistants that adapt to individual users’ preferences, backgrounds, and behavioral patterns  (Salemi et al., 2023; Zhang et al., 2024). These personalization systems leverage user-specific information to provide tailored responses, recommendations, and interactions. For example, LaMP (Large Language Models Meet Personalization) introduces comprehensive benchmarks for evaluating personalized text generation  (Salemi et al., 2023), while recent work explores personalized dialogue agents that maintain consistent user preferences across conversations  (Zhang et al., 2018).
@@
These developments showcase the remarkable ability of modern LLMs to facilitate and embody given personas, with applications spanning from personalized user assistance to sophisticated multi-agent collaborations in virtual environments  (Chan et al., 2023; Jiang et al., 2022).
@@
Recent advances in LLM-driven NPCs demonstrate significant potential for transforming player experiences. Cross-platform dialogue systems allow NPCs to maintain consistent interactions across both game environments and social platforms like Discord  (Song et al., 2024), creating unprecedented continuity in player-character relationships. Collaborative quest completion systems in Minecraft showcase how LLM-driven NPCs can work alongside human players to accomplish shared objectives  (Rao et al., 2024), while function-calling capabilities enable AI Game Masters or NPCs to manage complex game mechanics and narrative progression  (Chen et al., 2024). Furthermore, specialized datasets like MCPDial  (Yang et al., 2024) and PeaCoK  (Gao et al., 2023) provide rich persona-driven dialogue collections that enhance the authenticity and depth of NPC interactions, supporting the development of more sophisticated conversational agents in gaming environments.
The growing interest in persona-grounded gaming applications has culminated in organized competitions such as the Commonsense Persona-Grounded Dialogue Challenge (CPDC) 2025 (Sony AI, 2025). This competition invited submissions aimed at developing NPC agents that demonstrate both persona consistency and task execution capabilities in a fantasy Role-Playing Game (RPG) environment, as illustrated in Figure 1.
@@ 2.1 Agents for Game-Oriented Dialogue
Task-oriented systems are designed to complete specific tasks within larger projects, often as prerequisites for subsequent steps. This approach emphasizes efficient and effective task completion to achieve defined goals. The integration of agentic systems into such workflows enables agents to analyze problems, formulate plans, and execute them to achieve desired outcomes.
The performance of these agents, particularly in task-oriented dialogue (TOD), has been a key area of research. For instance,  (Kazi et al., 2024) established a benchmark for evaluating agent performance in TOD systems. They highlight that an agent’s planning capabilities can be assessed by focusing on the final stage of a simulation, the alignment of the proposed solution with the goal, or the methods the agent uses to interact with the system, such as gathering sufficient information to solve a problem.
@@ 4.1 API Track
In the API track, participants submit their work within specific environment and constraints such as the allowed LLM is GPT4o-mini (see D for full details). We focused on prompting methods. Our pipeline is illustrated in Figure 2. We systematically explored the following prompting approaches:
  • D (Deflanderization): Prompts the model to respond naturally and concisely while avoiding exaggerated role-playing.
@@ 4.2 GPU Track
We generated the datasets using gemini-2.5-pro-preview-05-06 (deepmind, 2025) for function-calling data and Gpt4o-mini (OpenAI, 2024) for dialogue data. The generated datasets consist of: Multi-turn (2,800 datapoints), Multi-turn reasoning (2,800 datapoints) for Task 2 (3.1.2) and Funtion-calling generation (328 data points) for Task 1 (3.1.1). Prompts used for data generation are provided in  C.1.
@@ 5.1 API Track
Adding few-shot examples (F) to the Deflanderization prompt provided clear benefits in Task 1 function name and argument matching (Table 2), with improvements of +0.092 and +0.133, respectively, on train.json. Notably, our best-performing API submission combined D-RW with two-turn few-shot examples, yielding the highest leaderboard placement (2nd on Task 3, 2nd on Task 1, and 5th on Task 2).
@@ 6 Discussion
Overall, our findings reveal complementary strategies across the API and GPU tracks. Prompting-based Deflanderization with few-shot grounding proved effective in low-resource API settings, while finetuned large models dominated the GPU track. Importantly, both tracks highlighted the challenge of balancing persona consistency with functional precision: methods that improved role-play fidelity sometimes hurt argument correctness, and vice versa. Future work should explore hybrid strategies that unify lightweight prompting with retrieval-augmented finetuning, enabling agents to sustain both accuracy and believability in fantasy RPG environments. See Our final ranking in Appendix F.
@@ Appendix F Final Leader Board
Table 4: our team Tu_Character_lab’s final result on API Track by AIcrowd Team. Task2 and Task3 also were evaluated by human while Task1 do only automatic evaluation.

### v2

@@ Abstract
Pasin Buakhaw1thanks: Equal contribution., Kun Kerdthaisong211footnotemark: 1, Phuree Phenhiran211footnotemark: 1,Pitikorn Khlaisamniang3, Supasate Vorathammathorn3,Piyalitt Ittichaiwong4,5thanks: Corresponding authors., Nutchanon Yongsatianchot [^2] footnotemark: 2 1Department of Computer Engineering and Digital Technology, Faculty of Engineering, Chulalongkorn University 2Faculty of Engineering, Thammasat School of Engineering, Thammasat University 3Artificial Intelligence Association of Thailand 4School of Biomedical Engineering & Imaging Sciences, King’s College London 5Siriraj Informatics and Data Innovation Center (SIData+), Faculty of Medicine, Siriraj Hospital, Mahidol University
@@ 1 Introduction
First, user persona-LLMs are designed as purpose-built assistants that adapt to individual users’ preferences, backgrounds, and behavioral patterns  (Salemi et al., 2023; li2402personalized). These personalization systems leverage user-specific information to provide tailored responses, recommendations, and interactions. For example, LaMP (Large Language Models Meet Personalization) introduces comprehensive benchmarks for evaluating personalized text generation  (Salemi et al., 2023), while another work explores personalized dialogue agents that maintain consistent user preferences across conversations  (Zhang et al., 2018).
@@
These developments showcase the remarkable ability of modern LLMs to facilitate and embody given personas, with applications spanning from personalized user assistance to sophisticated multi-agent collaborations in virtual environments  (wu2023large; Jiang et al., 2022).
@@
Recent advances in LLM-driven NPCs demonstrate significant potential for transforming player experiences. Cross-platform dialogue systems allow NPCs to maintain consistent interactions across both game environments and social platforms like Discord  (Song et al., 2024), creating unprecedented continuity in player-character relationships. Collaborative quest completion systems in Minecraft showcase how LLM-driven NPCs can work alongside human players to accomplish shared objectives  (Rao et al., 2024), while function-calling capabilities enable AI Game Masters or NPCs to manage complex game mechanics and narrative progression  (song2024you). Furthermore, specialized datasets like MCPDial  (alavi2024mcpdial) and PeaCoK  (Gao et al., 2023) provide rich persona-driven dialogue collections that enhance the authenticity and depth of NPC interactions, supporting the development of more sophisticated conversational agents in gaming environments.
The growing interest in persona-grounded gaming applications has culminated in organized initiatives such as the Commonsense Persona-Grounded Dialogue Challenge (CPDC) 2025 (Sony AI, 2025). This competition invited submissions aimed at developing NPC agents capable of demonstrating both persona consistency and task execution proficiency within a fantasy Role-Playing Game (RPG) environment, as illustrated in Figure 1.
Building upon this challenge, our work investigates multiple complementary strategies for enhancing the coherence and reliability of LLM-driven NPCs across diverse interaction settings. Specifically, we explore a Deflanderization prompting approach that mitigates character drift and preserves personality coherence during extended dialogues, ensuring balanced performance between dialogue generation and function execution. To further strengthen contextual grounding, we integrate a Retrieval-Augmented Generation (RAG) mechanism that incorporates memory and similarity-based retrieval from prior interactions, allowing NPCs to produce responses aligned with both in-game history and established world knowledge. Finally, we employ Supervised Finetuning (SFT) with Low-Rank Adaptation (LoRA) to refine model behavior at the parameter level, enhancing stylistic consistency and functional precision while remaining computationally efficient.
Together, these methods constitute a unified framework that examines the interplay between prompting, retrieval-augmented reasoning, and finetuned adaptation in achieving persona-consistent, context-aware, and goal-directed NPC performance within the CPDC 2025 setting.
@@
  • Explore the performance trade-offs between dialogue generation and function generation tasks using the proposed prompt engineering technique.
@@ 2.1 Agents for Game-Oriented Dialogue
Task-oriented systems are designed to efficiently complete specific tasks within larger workflows, often serving as prerequisites for later stages. Integrating agentic systems enhances these workflows by enabling agents to analyze problems, plan, and execute actions toward defined goals. Research on task-oriented dialogue (TOD) systems, such as  (Kazi et al., 2024), benchmarks agent performance by assessing planning effectiveness, goal alignment, and interaction methods used to gather information and achieve successful outcomes.
@@ 4.1 API Track
In the API track, participants submit their work within specific environment and constraints such as the allowed LLM is GPT-4o-mini (see D for full details). We focused on prompting methods. Our pipeline is illustrated in Figure 2. We systematically explored the following prompting approaches:
  • D (Deflanderization): Prompts the model to respond naturally and concisely while avoiding exaggerated role-playing. Our error analysis of the baseline setup from the challenge comparing generated responses with gold-standard outputs revealed that the baseline setup often produced overly elaborate and contextually diffuse outputs, focusing excessively on the narrative setting (e.g., adopting an RPG character persona) rather than addressing the immediate conversational intent to player. In contrast, the gold responses reflected a more human-like understanding of player requests and directly activated the appropriate functions with clarity.
@@ 4.2 GPU Track
We generated the datasets using gemini-2.5-pro-preview-05-06 (deepmind, 2025) for function-calling data and GPT-4o-mini (OpenAI, 2024) for dialogue data. The generated datasets consist of: Multi-turn (2,800 data points), Multi-turn reasoning (2,800 data points) for Task 2 (3.1.2) and Funtion-calling generation (328 data points) for Task 1 (3.1.1). Prompts used for data generation are provided in  C.1.
@@ 6 Discussion
Overall, our findings reveal complementary strategies across the API and GPU tracks. Prompting-based Deflanderization with few-shot grounding proved effective in low-resource API settings, while finetuned large models dominated the GPU track. Importantly, both tracks highlighted the challenge of balancing persona consistency with functional precision: methods that improved role-play fidelity sometimes hurt argument correctness, and vice versa. Future work should explore hybrid strategies that unify lightweight prompting with retrieval-augmented finetuning, enabling agents to sustain both accuracy and believability in fantasy RPG environments. Our final rankings are in Appendix F.
Acknowledgments
This research was supported by the Faculty of Engineering, Thammasat School of Engineering, Thammasat University also thanks to PreceptorAI that provides API for generate additional training data.
@@ Appendix F Final Leader Board
Table 4: our team Tu_Character_lab’s final result on API Track by AIcrowd Team. Task 2 and Task 3 also were evaluated by human while Task 1 was evaluated automatically.